Circular No. 6169

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

PKS 2155–304 W. T. Vestrand and J. G. Stacy, University of New Hampshire; and

P. Sreekumar, Goddard Space Flight Center, NASA, report: “Our analysis of data taken during 1994 Nov. 15–29 by the Energetic Gamma Ray Experiment Telescope (EGRET) aboard the Compton Gamma Ray Observatory has resulted in the first detection of high-energy γ-ray emission from the prototypical, x-ray-selected BL Lac object PKS 2155–304. The photon power-law index derived for the 30-MeV to 10-GeV emission has an index of −1.69 ± 0.26 and an integral flux above 100 MeV of (2.7 ± 0.7) × 10^{−7} photon cm^{−2}. The spectral hardness of the source and its proximity (z = 0.117) make it an ideal candidate for measurements at TeV γ-ray energies. Since a detection of this object at TeV energies will place important constraints on intergalactic infrared radiation fields, we urge observers of TeV γ-ray sources to search for emission from PKS 2155–304.”

GRAVITATIONAL MICROLENSING EVENTS IN PROGRESS The MACHO collaboration (cf. IAUC 6155, plus D. Welch) reports

the probable discovery of two more gravitational microlensing events in progress toward the galactic bulge: “The stars in question are events 95-7, α = 18^{h}13^{m}29^{s}.3, δ = −26^{o}13′58′′ (equinox 2000.0); and 95-8, 18^{h}16^{m}46^{s}.0, −26^{o}11′43′′. These stars were constant in brightness during 1993, at R ∼ 17.5 (V ∼ 18.2) and 16.3 (17.1), respectively, and both have brightened by 0.9 mag as of 1995 May 3. Microlensing fits to data obtained with the Mt. Stromlo 1.3-m reflector indicate that event 95-7 is currently near its peak amplification, while event 95-8 should reach its peak between May 7 and 13. Accurate photometry near the peak of event 95-8 may yield valuable information about the nature of the microlensing object. The MACHO collaboration will henceforth maintain an online catalogue of ongoing candidate microlensing events (in lieu of announcing via IAUC) for each detected candidate event. The MACHO collaboration is also working to establish a world-wide network to provide continuous photometry of microlensing events to enhance the probability of detecting departures from simple microlensing, in particular those indicating the existence of extrasolar planets. Further information is available via anonymous ftp from

darkstar.astro.washington.edu in the macho/Alert directory, and at
the WWW site http://darkstar.astro.washington.edu/. Questions or

comments should be directed to macho@astro.washington.edu.”

1995 May 5 Daniel W. E. Green